# Energy-Enhanced Expansion of the Standard Model Effective Field Theory

### v1

@@ V.3 $\bar{q}q\rightarrow jjHH$
Carrying out the $\lambda$ power counting, we find that the dominant SM contribution (at large energy) is the rightmost topology b.) of Fig 3, scaling as $\lambda^{4}$. Without additional assumptions, the dominant SMEFT contribution comes at $\mathcal{O}(\lambda^{6})$ from ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$, which modify the $H^{2}V^{2}$ coupling. If we zero these operators to take into account the fact that they are bounded from $H\to\gamma\gamma$ measurements, while also zeroing operators either bounded by LEP (such as $c_{\psi^{2}H^{2}D}$ ) or with non-SM helicity structure, the first SMEFT effects to enter are $\mathcal{O}(\lambda^{8})$ and come from operator classes $c_{H^{4}D^{2}}$ at dimension six and $c_{\psi^{2}H^{2}D^{3}},c_{\psi^{2}H^{2}XD},c_{\psi^{4}H^{2}}$ at dimension eight. The impact of two of these three ($c_{\psi^{2}H^{2}D^{3}}$ and $c_{\psi^{4}H^{2}}$) on high-energy single Higgs VBF ($pp\to jjH$) production was recently explored in Ref. Assi:2024zap . These operators were carefully studied in that work and confirmed to be the most significant contributors, thereby validating our prescription.
@@ VI Discussion
Several caveats are in order. First, throughout this work we assume that new physics is weakly coupled—that is, all Wilson coefficients $c_{i}/\Lambda$ are taken to be of comparable magnitude (apart from the standard loop suppression). Second, our power counting is based on on-shell conditions, so vertices that vanish by the equations of motion are not properly ordered within this scheme. Third, while our analysis is carried out to leading order in $\lambda$, sub-leading corrections may be relevant for certain observables and warrant further study. Finally, our illustrative assumptions—such as setting to zero those operators whose interference with the SM is suppressed—are chosen to simplify the discussion; a more complete phenomenological analysis would need to account for additional effects from flavor structure, loop-level matching, and off-shell dynamics.
@@ Acknowledgements
We thank Kelci Mohrman and Pavel Nadolsky for their helpful feedback. The work of AM is partially supported by the National Science Foundation under Grant Numbers PHY–2112540 and PHY–2412701. The work of BA is supported by the DOE grant number DE–SC0011784 and NSF grant number OAC–2103889 as well as the Fermi National Accelerator Laboratory (Fermilab). Fermilab is managed by Fermi Research Alliance, LLC (FRA), acting under Contract No. DE–AC02–07CH11359. This work was performed in part at the Aspen Center for Physics, with support for BA by a grant from the Simons Foundation (1161654,Troyer).

### v2

@@ V.3 $\bar{q}q\rightarrow jjHH$
Carrying out the $\lambda$ power counting, we find that the dominant SM contribution (at large energy) is the rightmost topology b.) of Fig 3, scaling as $\lambda^{4}$. Without additional assumptions, the dominant SMEFT contribution comes at $\mathcal{O}(\lambda^{6})$ from $\bm{c_{H^{2}X^{2}}}$, which modify the $H^{2}V^{2}$ coupling. If we zero these operators to take into account the fact that they are bounded from $H\to\gamma\gamma$ measurements, while also zeroing operators either bounded by LEP (such as $c_{\psi^{2}H^{2}D}$ ) or with non-SM helicity structure, the first SMEFT effects to enter are $\mathcal{O}(\lambda^{8})$ and come from operator classes $c_{H^{4}D^{2}}$ at dimension six and $c_{\psi^{2}H^{2}D^{3}},c_{\psi^{2}H^{2}XD},c_{\psi^{4}H^{2}}$ at dimension eight. The impact of two of these three ($c_{\psi^{2}H^{2}D^{3}}$ and $c_{\psi^{4}H^{2}}$) on high-energy single Higgs VBF ($pp\to jjH$) production was recently explored in Ref. Assi:2024zap . These operators were carefully studied in that work and confirmed to be the most significant contributors, thereby validating our prescription.
@@ VI Discussion
Several caveats are in order. First, throughout this work we assume that new physics is weakly coupled—that is, all Wilson coefficients $c_{i}/\Lambda$ are taken to be of comparable magnitude (apart from the standard loop suppression). Second, our power counting is based on on-shell conditions, so vertices that vanish by the equations of motion are not properly ordered within this scheme. Third, while our analysis is carried out to leading order in $\lambda$, sub-leading corrections may be relevant for certain observables and warrant further study. Thus, the $\lambda$ counting we advocate for is a guide; large numerical or coupling prefactors that appear in explicit calculations can override it, so final priorities on what orders in $\lambda$ to keep should be set after computing the full amplitude. Finally, our illustrative assumptions—such as setting to zero those operators whose interference with the SM is suppressed—are chosen to simplify the discussion; a more complete phenomenological analysis would need to account for additional effects from flavor structure, loop-level matching, and off-shell dynamics.
@@ Acknowledgements
We thank Kelci Mohrman, Pavel Nadolsky and Nick Smith for their helpful feedback. The work of AM is partially supported by the National Science Foundation under Grant Numbers PHY–2112540 and PHY–2412701. The work of BA is supported by the DOE grant number DE–SC0011784 and NSF grant number OAC–2103889 as well as the Fermi National Accelerator Laboratory (Fermilab). This manuscript has been authored by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work was performed in part at the Aspen Center for Physics, with support for BA by a grant from the Simons Foundation (1161654,Troyer).